OpenStudy (anonymous):

Simplify: 18 – 4 + 12 divided 2 I keep getting 8, but it's not one of the choices. Please explain the steps

OpenStudy (notamathgenius):

Order of operations if i am not mistaken you got PEMDAS add first then subtract then divide by 2

OpenStudy (anonymous):

So the answer would be 1? I thought for PEMDAS you Divide first, add, then subtract.

OpenStudy (notamathgenius):

You cant divide until you have an answer is 1 an answer? I got 7

OpenStudy (anonymous):

Yeah, 1 is one of the choices. 7 is not one of the choices.

OpenStudy (notamathgenius):

Ok I would go with your answer i must have dont it backwards, let me re do it

OpenStudy (notamathgenius):

yea i got 1

OpenStudy (anonymous):

18 – 4 + 12=26 then 26/2=13

OpenStudy (anonymous):

(18-4+12/2) or 18-4+12/2 ?

OpenStudy (notamathgenius):

Its order of operations so addition has to be put first

OpenStudy (anonymous):

18 – 4 + 12 / 2

OpenStudy (anonymous):

Wouldnt you use PEMDAS?

OpenStudy (anonymous):

is it like this\[\frac{ 18-4+12 }{ 2}\]

OpenStudy (anonymous):

So the answer is 13? For some reason I thought you divide first

OpenStudy (anonymous):

yes it is 13

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

13 is incorrect :(

OpenStudy (notamathgenius):

did you try 1

OpenStudy (anonymous):

It won't let me do it again, but i'm pretty sure

OpenStudy (notamathgenius):

ahhh

OpenStudy (notamathgenius):

You're welcome

OpenStudy (anonymous):

your problem is not correct that is why u did not get the answer

OpenStudy (anonymous):

i guess it is like this \[\frac{ 18-4-12 }{ 2}=1\]

OpenStudy (anonymous):

My problem is correct though, this is the question Simplify: 18 – 4 + 12/ 2

OpenStudy (anonymous):

Oh ok! So it should be 1.

OpenStudy (anonymous):

are u sure it is 1

OpenStudy (anonymous):

No, I thought that's what you were saying? I think this problem is just written bad

OpenStudy (anonymous):

well, i am sure about my solution if it is like the way i wrote

OpenStudy (anonymous):

I'm going to ask my teacher. Thanks anyways